OpenStudy (anonymous):

What is the multiplicative identity of the complex number -3+5i? I'm having a hard time understanding multiplicative identities and this question is on my assignment. Please help!

myininaya (myininaya):

the multiplicative identity it 1 since for any a in the complex number we have a*1=1*a=a

myininaya (myininaya):

If however you mean multiplicative inverse of a then you need a number such that when you multiply it to a you get 1.

myininaya (myininaya):

for example the multiplicative inverse of 2 is 1/2 because 2*(1/2)=1

OpenStudy (anonymous):

your wrong

myininaya (myininaya):

if you wanted the multiplicative inverse of -3+5i then you would need to figure out for what complex number z you have (-3+5i)z=1

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The answer is abviously simple if u just do all the math and put all the clues together the pie represents about a week ago. see(-3+5i)z= 3.249

myininaya (myininaya):

anyways z is a complex number z=a+bi just multiply (-3+5i) and (a+bi) as a first step if indeed it is the multiplicative inverse you seek

myininaya (myininaya):

then you will compare both sides to find out what a and b should be

OpenStudy (anonymous):

nononononononononononnonononononononoononononononoonoooooo

myininaya (myininaya):

for example say we wanted to find the multiplicative inverse of 2-5i then say it is a+bi so we have that (2-5i)(a+bi)=1 multiply the left hand side \[2a+2bi-5ia-5bi^2=1 \\ 2a+2bi-5ai-5b(-1)=1 \\ 2a+(2b-5a)i+5b=1 \\ (2a+5b)+(2b-5a)i=1 \\ \text{ you have two equations to solve } \\ 2a+5b=1 \text{ and } 2b-5a=0 \]

myininaya (myininaya):

so you have 2a+5b=1 and 2b=5a => b=5a/2 plug that into 2a+5b=1 so you have 2a+5(5a/2)=1 therefore 2a+25a/2=2 => 4a+25a=2 a=2/29 so b=5a/2=5/2* 2/29=5/29

myininaya (myininaya):

so the multiplicative inverse of (2-5i) is (2/29+5i/29)

OpenStudy (anonymous):

Thank you both! You have given me much to think about.